How many transitive relatuon are there ona set with m elements

If m=1 , m=2 and m=3

Solve the question


1
Expert's answer
2020-08-26T15:55:26-0400

A relation "R" on a set "A" is transitive if "(a, b)\in R, (b, c)\in R" implies "(a, c)\in R."

a) "m=1"

Let "a" be the only element in "A." All possible relations are:

"(a, a)" Transitive

"\text{\o}" Transitive

There are two transitive relations on the set "A."


b) "m=2"

Let "a, b" be the elements in "A." Then there are 4 possible ordered pairs on the set "A"


"(a, a), (a, b), (b, a), (b, b)"

There are 16 relations in all. The only way a relation can fail to be transitive is to contain both "(a, b)" and "(b, a)". If the relation contains all four elements it is transitive.

Hence there are 13 transitive relations:

"\text{\o}, \{(a,a)\},\{(a,b)\},\{(b,a)\}, \{(b,b)\},"

"\{(a,a), (a,b)\},\{(a,a), (b,a)\},\{(a,a), (b,b)\},"

"\{(a,b), (b,b)\},\{(b,a), (b,b)\},"

"\{(a,a), (a,b),(b,b)\},\{(a,a), (b,a),(b,b)\},"

"\{(a,a), (a,b),(b,a)(b,b)\}"

There are 13 transitive relations on the set "A."


c) "m=3"

Let "a, b" and "c" be the elements in "A." Then there are 9 possible ordered pairs on the set "A"


"(a, a), (a, b), (a,c), (b, a), (b, b),(b,c),(c,a),(c,b),(c,c)"


For each of the ordered 9 pairs, we have 2 choices: the element is in the set or the element is not in the set. By the product rule


"2\cdot 2\cdot 2\cdot...\cdot2=2^9=512"

There are 512 relations in all.

From https://en.wikipedia.org/wiki/Transitive_relation#Counting_transitive_relations

No general formula that counts the number of transitive relations on a finite set is known.

Using technology to generate all possible relations, we would then obtain 171 transitive relations among the 512 relations.